How many quarts of milk containing 4% of butterfat and how many quarts of cream containing 24% of butterfat must be mixed to make 200 quarts of milk containing 12% of butterfat?
Answer by scott8148(5891)

let x=milk, so 200-x=cream ___ x(4%)+(200-x)(24%)=200(12%)

.04x+48-.24x=24 ___ 24=.20x